What is the value of sin(\thetaθ) for a setup using a 709 micron/micrometer slit with a minima of order 9 has a wavelength of 3080 nm?
Answer #1

setup is 713 micron/micrometer

=> distabce between slit , d = 1/(709) = 1.410*10^-3 m

accordint to the slit equation

d*sin(theta) = m*wavelength

=> sin(theta) = 9*3080*10^-9/(1.410*10^-3) =0.0196